OpenStudy (anonymous):

How many grams of NaOH are required to neutralize 25.0 mL of 0.50 M HCl?

OpenStudy (anonymous):

2gram of NAOH is required

OpenStudy (anonymous):

\[? gr NaOH = 25.0 ml HCl (\frac{ 0.5 mol HCl }{1000 ml HCl })(\frac{ 1mol NaOH }{1 mol HCl })(\frac{ 40.0gr NaOH }{ 1mol NaOH }) = 0.5 gr NaOH\]